# Feature-based prediction of properties of cross-linked epoxy polymers by molecular dynamics and machine learning techniques

### v1

@@ 5 Conclusions
In this study, we used molecular dynamics simulations and machine learning techniques to get in depth understanding on the structural features and process parameters affecting the mechanical properties of epoxy polymers. We used MD cross-linking simulations to develop EPON-862 based, epoxy polymeric systems with different (11) types of curing agents (hardeners) in addition to the conventionally used ones. We determined the mechanical properties (elastic modulus and yield strength) of these cross-linked polymeric systems from the MD strain simulations and validated with those from experimental studies reported in the literature. While the elastic modulus determined from MD simulations (using different strain rates) was found to be within the range of values reported in literature, the yield strength was found to be higher by a factor of 10 due to very high strain rate adopted in MD simulations as against the strain rate conventionally used during experiments since the strength is more sensitive to strain rate. From these simulations, we determined the influence of different parameters like the stoichiometric ratio, curing percentage and the type of hardener on the mechanical properties of epoxy polymers. We classified the hardeners into different categories like the ones with aromatic rings, aliphatic rings and aliphatic chains and the stress-strain response of each category of hardeners was analysed individually. From the analysis, it has been identified that, in the case of hardeners with aromatic rings, the hardener (DDM) with two aromatic rings possess less mechanical properties than the ones with single aromatic ring and in the case of hardeners with aliphatic chains and aliphatic rings, the hardener with more reactive amine groups possess higher mechanical properties. We further used GPR based ML model using which the mechanical properties of epoxy polymers can be predicted from their structural features extracted from cheminformatics tool. We could also determine the most important structural features whose information may be required for predicting the elastic modulus and yield strength accurately using sequential backward selection technique. We identified that parameters like cross-link percentage, number of carbon atoms, and primary amines are most important for predicting the elastic modulus while parameters like cross-link percentage, number of aliphatic rings, primary amines and secondary amines are most important for predicting the yield strength. The developed model performs with an accuracy (MAE) of 0.11510.0176 GPa for predicting elastic modulus and 0.0054 0.0011 GPa for predicting yield strength (note to factor the strength by 10 for comparing with conventional experiments). The model also demonstrates the Score of 0.720.07 and 0.890.05 while predicting elastic modulus and yield strength respectively. The model has also found to perform best in comparison to conventional regression models for predicting yield strength whereas for predicting elastic modulus, it performed slightly lower than ridge regression. The proposed method can thus be extended to simulate a larger number of epoxy polymeric systems virtually and determine their properties in an efficient way. This will further lead towards increased performance space of the epoxy polymers and also aid in designing the polymers by merely selecting the constituents with appropriate structural features.

### v2

@@ 5 Conclusions
In this study, we used molecular dynamics simulations and machine learning techniques to get in depth understanding on the structural features and process parameters affecting the mechanical properties of epoxy polymers. We used MD cross-linking simulations to develop EPON-862 based, epoxy polymeric systems with different (11) types of curing agents (hardeners) in addition to the conventionally used ones. We determined the mechanical properties (elastic modulus and yield strength) of these cross-linked polymeric systems from the MD strain simulations and validated with those from studies reported in the literature. The elastic modulus determined from MD simulations (using different strain rates) was found to be within the range of values (experimental and computational) reported in literature. The yield strength determined from MD simulations was also found to be within the range of values reported from MD simulations in the literature. However, the yield strength determined from MD simulations is found to be 3 times higher than the values reported from experimental investigations due to very high strain rate adopted in MD simulations as against the strain rate conventionally used during experiments since the strength is more sensitive to strain rate. From these simulations, we determined the influence of different parameters like the stoichiometric ratio, curing percentage and the type of hardener on the mechanical properties of epoxy polymers. We classified the hardeners into different categories like the ones with aromatic rings, aliphatic rings and aliphatic chains and the stress-strain response of each category of hardeners was analysed individually. From the analysis, it has been identified that, in the case of hardeners with aromatic rings, the hardener (DDM) with two aromatic rings possess less mechanical properties than the ones with single aromatic ring and in the case of hardeners with aliphatic chains and aliphatic rings, the hardener with more reactive amine groups possess higher mechanical properties. We further used GPR based ML model using which the mechanical properties of epoxy polymers can be predicted from their structural features extracted from cheminformatics tool. We could also determine the most important structural features whose information may be required for predicting the elastic modulus and yield strength accurately using sequential backward selection technique. We identified that parameters like molecular weight, fraction of carbon atoms that are SP3 hybridized, total number of carbon atoms, primary amines and secondary amines are most important for predicting the elastic modulus while parameters like molecular weight, stoichiometric ratio between resin and hardener, curing percentage, total number of oxygen atoms, aromatic rings and primary amines are most important for predicting the yield strength. The developed model performs with an accuracy (MAE) of 0.11570.0190 GPa for predicting elastic modulus and 0.0067 0.0016 GPa for predicting yield strength. The model also demonstrates the Score of 0.720.05 and 0.810.02 while predicting elastic modulus and yield strength respectively. The model has also found to perform best in comparison to conventional regression models for predicting yield strength whereas for predicting elastic modulus, it performed slightly lower than gradient boosting. The proposed method can thus be extended to simulate a larger number of epoxy polymeric systems virtually and determine their properties in an efficient way. This will further lead towards increased performance space of the epoxy polymers and also aid in designing the polymers by merely selecting the constituents with appropriate structural features.